A 5kg ball is kicked into the air with a speed of 14m/s after 1.42seconds, it reaches the peak of its trajectory calculate the time the change in position The initial velocity the final velocity the average velocity and the acceleration of the object on its way up(before it starts to fall back down) use SI units


1
Expert's answer
2021-03-26T11:37:58-0400

"t=1.42~s,"

"v_1=v=14~\frac ms,"

"v_2=0~\frac ms,"

"h=\frac{v^2}{2g},"

"<v>=\frac ht=\frac{v^2}{2gt}=\frac{14^2}{2\cdot 9.8\cdot 1.42}=7~\frac ms,"

"a=-g=-9.8~\frac{m}{s^2}."